# Tate modules as condensed modules

### v1

@@ Abstract.
We prove that the category of countable Tate modules over an arbitrary discrete ring embeds fully faithfully into that of condensed modules. If the base ring is of finite type, we characterize the essential image as generated by the free module of infinite countable rank under direct sums, duals and retracts. In the $\infty$-categorical context, we establish a fully faithful embedding of the $\infty$-category of countable Tate objects in perfect complexes, with uniformly bounded tor-amplitude, into the derived $\infty$-category of condensed modules. The boundedness assumption is necessary to ensure fullness, as we prove via an explicit counterexample in the unbounded case.
@@ Introduction
More precisely, we prove in Section 1 that for any commutative ring $R$ there is a fully faithful exact functor
@@
Section 3 is devoted to extend the fully faithfulness result to the $\infty$-category of Tate $R$-modules (as defined in [Hen17b]). This extension does not seem to be a trivial consequence of the 1-categorical result; in fact our approach to the $\infty$-categorical statement will be completely independent of the first two sections, avoiding the need for topological modules. Actually this $\infty$-categorical result turns out to hold only if we limit ourselves to perfect complexes (replacing finitely generated projective modules in the $\infty$-categorical setting) which are of fixed tor-amplitude. The result is achieved in Corollary 3.3 as a consequence of a more general embedding (see Theorem 3.2)
$\mathrm{Pro}_{\aleph_{0}}^{[a,b]}\mathopen{}\mathclose{{}\left(\mathscr{D}(R)}\right)\lhook\joinrel\relbar\joinrel\rightarrow\mathscr{D}(\mathsf{Cond}_{R})$
of Pro-complexes (which contain Tate objects in perfect complexes) with tor-amplitude in $[a,b]$ into the derived $\infty$-category $\mathscr{D}(\mathsf{Cond}_{R})$ of condensed $R$-modules. Finally, an interesting counterexample to full faithfulness for arbitrary unbounded perfect complexes is given in Section 4.
@@
Acknowledgments. We started this project together with Andrea Maffei, and he contributed some important ideas; in particular, Section 1 should be considered as joint work with him. A first version of Proposition 1.5 was obtained by Francesco Iacca in his Master Thesis [Iac24]. We thank Chris Brav, Benjamin Hennion, Francesco Iacca for extremely useful discussions on the topics of this paper.
@@ 3. $\infty$-categorical embedding in the bounded case
Let $R$ be an arbitrary ring. In this section, we show that countable pro-objects in the derived $\infty$-category of a ring $R$, if uniformly bounded in the sense of tor-amplitude, embed fully faithfully into the derived $\infty$-category of condensed $R$-modules.
@@
  • Following [Hen17b, Definition 2.1], we let $\mathscr{T}\mathrm{ate}_{R}:=\mathscr{T}\mathrm{ate}(\mathrm{Perf}_{R})$ be the stable $\infty$-category of Tate objects in perfect complexes of $R$-modules. In this section, we will restrict our attention to its full subcategory $\mathscr{T}\mathrm{ate}_{\aleph_{0},R}$ of countable Tate objects in $\mathrm{Perf}_{R}$, by which we mean those Tate objects that can be expressed using only countable filtered and cofiltered diagrams.
  • Given integers $a$ and $b$, let $\mathscr{D}(R)^{[a,b]}$ denote the full subcategory of the derived $\infty$-category $\mathscr{D}(R)$ of $R$-modules consisting of objects with tor-amplitude in $[a,b]$, where we use cohomological grading conventions.
  • Let $\mathrm{Pro}^{[a,b]}\mathopen{}\mathclose{{}\left(\mathscr{D}(R)}\right)$ denote the full subcategory of the $\infty$-category $\mathrm{Pro}\mathopen{}\mathclose{{}\left(\mathscr{D}(R)}\right)$ consisting of Pro-objects that are uniformly of tor-amplitude in $[a,b]$, i.e. that can be written as a cofiltered diagrams taking values in $\mathscr{D}(R)^{[a,b]}$. The intersection of this $\infty$-category with $\mathscr{T}\mathrm{ate}_{\aleph_{0},R}$ will be denoted $\mathscr{T}\mathrm{ate}_{\aleph_{0},R}^{[a,b]}$.
  • Given an $\infty$-category $\mathscr{C}$ with finite limits, we let $\mathsf{Cond}(\mathscr{C})$ be the $\infty$-category of condensed objects in $\mathscr{C}$, defined as accessible $\mathscr{C}$-valued sheaves on profinite sets. For $\mathscr{C}=\mathscr{D}(R)$, note that there is an equivalence $\mathsf{Cond}(\mathscr{D}(R))\simeq\mathscr{D}(\mathsf{Cond}_{R})$.
Observe that the fully faithful functor $\mathsf{Mod}_{R}\hookrightarrow\mathsf{Cond}_{R}$ defined in (1.4) induces a fully faithful functor of stable $\infty$-categories $\mathscr{D}(R)\hookrightarrow\mathscr{D}(\mathsf{Cond}_{R})$, which is left adjoint to evaluation at $*\in\mathsf{ProFin}$. Beware that, as opposed to the previous sections, we will not distinguish between an object in $\mathscr{D}(R)$ and its image through this functor.
@@
For every integers $a,b\in\mathbb{Z}$, the restriction
$\mathrm{Pro}_{\aleph_{0}}^{[a,b]}\mathopen{}\mathclose{{}\left(\mathscr{D}(R)}\right)\longrightarrow\mathscr{D}(\mathsf{Cond}_{R})$
of the canonical functor (3.1) is fully faithful.
Corollary 3.3.
The restriction $\mathscr{T}\mathrm{ate}_{\aleph_{0}}^{[a,b]}\mathopen{}\mathclose{{}\left(\mathrm{Perf}_{R}}\right)\longrightarrow\mathscr{D}(\mathsf{Cond}_{R})$ of the functor (3.1) to Tate objects is fully faithful.
Proof of theorem 3.2.
Let $V\colon I\to\mathscr{D}(R)$ and $W\colon J\to\mathscr{D}(R)$ be countable pro-objects in $\mathscr{D}(R)$ with tor-amplitude in $[a,b]$. Using the description of the mapping spaces in pro-categories, we need to show that the canonical map
$\lim_{j\in J}\operatorname*{colim}_{i\in I^{\mathrm{op}}}\operatorname{Map}_{\mathscr{D}(R)}(V_{i},W_{j})\longrightarrow\operatorname{Map}_{\mathscr{D}(\mathsf{Cond}_{R})}\big{(}\lim_{i\in I}V_{i},\lim_{j\in J}W_{j}\big{)}$ (3.2)
is an equivalence, where the limits in the right hand side are taken in $\mathscr{D}(\mathsf{Cond}_{R})$. Using the universal property of limits, the desired result will be a consequence of the following proposition. ∎
Proposition 3.4.
Let $V\colon I\to\mathscr{D}(\mathsf{Cond}_{R})$ be a countable pro-condensed object and $W\in\mathscr{D}(R)$ be a discrete module. Assume that the following boundedness hypothesis is satisfied:
there exists $a,b\in\mathbb{Z}$ such that either $V$ factors through $\mathscr{D}(\mathsf{Cond}_{R})^{[a,b]}$ or $W$ is in $\mathscr{D}(R)^{[a,b]}$. ($\star$)
Then the canonical morphism
$\gamma\colon\operatorname*{colim}_{i\in I^{\mathrm{op}}}\operatorname{Map}_{\mathscr{D}(\mathsf{Cond}_{R})}(V_{i},W)\longrightarrow\operatorname{Map}_{\mathscr{D}(\mathsf{Cond}_{R})}\big{(}\lim_{i\in I}V_{i},W\big{)}$ (3.3)
@@
The rest of the section in devoted to the proof of proposition 3.4. We will proceed in two steps, proving that $\gamma$ is both a monomorphism of $\infty$-groupoids (i.e. an inclusion of connected components) and an effective epimorphism (i.e. surjective on $\pi_{0}$); these two statements correspond respectively to lemmas 3.6 and 3.8.
Notation 3.5.
Given an index $i\in I$, let $p_{i}$ denote the projection $\lim\limits_{j\in I}V_{j}\to V_{i}$ and $\iota_{i}\colon\mathrm{fib}(p_{i})\to\lim\limits_{j\in I}V_{j}$ its fiber.
Lemma 3.6.
The morphism $\gamma$ from (3.3) is a monomorphism.
@@
We wish to prove that the non-empty fibers of $\gamma$ are contractible. Consider a morphism $f\colon V_{i_{0}}\to W$ for some $i_{0}\in I$ and let $Z$ denote the fiber of $\gamma$ at the point $\gamma(f)\simeq f\circ p_{i_{0}}$, with its natural morphism $\jmath\colon Z\to\operatorname*{colim}_{j}\operatorname{Map}(V_{j},W)$. Let $\sigma\colon K\to Z$ be a morphism from a compact $\infty$-groupoid $K\in\mathscr{S}^{\omega}$. To prove the lemma, it is sufficient to show that $\sigma$ is nullhomotopic. Since $K$ is a compact object in $\mathscr{S}$ and $I$ is filtered, the morphism $\jmath\circ\sigma\colon K\to\operatorname*{colim}_{j}\operatorname{Map}(V_{j},W)$ must factor through $\operatorname{Map}(V_{i},W)$ for some index $i\in I$. Choosing $i$ large enough, we may assume that there is a map $i_{0}\to i$ in the poset $I$, so that we have an equivalence $\gamma(f)\simeq f\circ p_{i}$. Therefore the restriction of $\gamma(f)$ along the canonical map $\iota_{i}\colon\mathrm{fib}(p_{i})\to\lim V$ is the zero morphism and thus we obtain a lift in the diagram
[…]
using that the right most square is cartesian. This implies that $\sigma$ is nullhomotopic, as desired. ∎

### v2

@@ Abstract.
We prove that the category of countable Tate modules over an arbitrary discrete ring embeds fully faithfully into that of condensed modules. If the base ring is of finite type, we characterize the essential image as generated by the free module of infinite countable rank under direct sums, duals and retracts. In the $\infty$-categorical context, under the same assumption on the base ring, we establish a fully faithful embedding of the $\infty$-category of countable Tate objects in perfect complexes, with uniformly bounded tor-amplitude, into the derived $\infty$-category of condensed modules. The boundedness assumption is necessary to ensure fullness, as we prove via an explicit counterexample in the unbounded case.
@@ Introduction
Outline of the paper. We prove in Section 1 that for any commutative ring $R$ there is a fully faithful exact functor
@@
Section 3 is devoted to extend the fully faithfulness result to the $\infty$-category of Tate $R$-modules, as defined in [Hen17b]. Contrary to the 1-categorical embedding, but as in Section 2, we will need to assume the base ring $R$ to be of finite type over $\mathbb{Z}$. The main result is achieved in Theorem 3.2 and asserts that the canonical realization functor
$\mathscr{T}\mathrm{ate}_{\aleph_{0},R}^{\mathrm{b}}\longrightarrow\mathscr{D}(\mathsf{Cond}_{R})$
from bounded countable Tate objects in perfect complexes to the derived $\infty$-category $\mathscr{D}(\mathsf{Cond}_{R})$ of condensed $R$-modules is fully faithful.
Restricting to Tate objects that are bounded (in the sense of having uniformly finite tor-amplitude, see Definition 3.1) is necessary, as we demonstrate in Section 4 by providing an explicit counterexample to full faithfulness for the $\infty$-category $\mathscr{T}\mathrm{ate}_{{\aleph_{0}}}(\mathrm{Perf}_{R})$.
After a first version of this paper was released, P. Scholze kindly pointed out a mistake in the $\infty$-categorical full faithfulness: the property of $R$ being a finitely generated $\mathbb{Z}$-algebra is necessary (see Remark 3.5). He also made us aware of the recent paper [BL24]. There is some overlap between [BL24, Section 2.2] and our results in Section 3, see Remark 3.4 for more details.
@@
Acknowledgments. We started this project together with Andrea Maffei, and he contributed some important ideas; in particular, Section 1 should be considered as joint work with him. A first version of Proposition 1.5 was obtained by Francesco Iacca in his Master Thesis [Iac24]. We thank Chris Brav, Benjamin Hennion, Francesco Iacca for extremely useful discussions on the topics of this paper. We thank Peter Scholze for pointing out an error in the first version of this paper and for making us aware of the paper [BL24].
@@ 3. $\infty$-categorical embedding in the bounded case
Let $R$ be a ring of finite type over $\mathbb{Z}$. In this section, we show that countable Tate objects in perfect complexes of $R$-modules, if uniformly bounded in the sense of tor-amplitude, embed fully faithfully into the derived $\infty$-category of condensed $R$-modules.
@@
  • Following [Hen17b, Definition 2.1], we let $\mathscr{T}\mathrm{ate}_{R}:=\mathscr{T}\mathrm{ate}(\mathrm{Perf}_{R})$ be the $\infty$-category of Tate objects in perfect complexes of $R$-modules. Recall that this is the smallest stable idempotent complete subcategory of $\mathrm{Pro}\,\mathrm{Ind}(\mathrm{Perf}_{R})$ containing both the essential images of $\mathrm{Pro}(\mathrm{Perf}_{R})$ and $\mathrm{Ind}(\mathrm{Perf}_{R})\simeq\mathscr{D}(R)$.
  • Restricting to countable cofiltered diagrams gives the full subcategory $\mathrm{Pro}_{\aleph_{0}}(-)\subset\mathrm{Pro}(-)$ of countable pro-objects; and dually for $\mathrm{Ind}_{\aleph_{0}}(-)\subset\mathrm{Ind}(-)$. We define the $\infty$-category $\mathscr{T}\mathrm{ate}_{\aleph_{0},R}$ of countable Tate objects in $\mathrm{Perf}_{R}$ as the intersection of $\mathscr{T}\mathrm{ate}_{R}$ with $\mathrm{Pro}_{\aleph_{0}}(\mathrm{Ind}_{\aleph_{0}}(\mathrm{Perf}_{R}))$.
  • Given integers $a$ and $b$, let $\mathrm{Perf}_{R}^{[a,b]}$ denote the full subcategory of the stable $\infty$-category $\mathrm{Perf}_{R}$ of perfect complexes of $R$-modules consisting of objects with tor-amplitude in $[a,b]$, where we use cohomological grading conventions. We may form the $\infty$-category $\mathrm{Pro}_{\aleph_{0}}\mathrm{Ind}_{\aleph_{0}}\mathopen{}\mathclose{{}\left(\mathrm{Perf}_{R}^{[a,b]}}\right)$ of countable pro-ind-perfect complexes that are uniformly of tor-amplitude in $[a,b]$. The intersection of this $\infty$-category with $\mathscr{T}\mathrm{ate}_{\aleph_{0},R}$ will be denoted $\mathscr{T}\mathrm{ate}_{\aleph_{0},R}^{[a,b]}$.
  • Taking the union of these subcategories $\mathscr{T}\mathrm{ate}_{\aleph_{0},R}^{[a,b]}$ for all $a,b\in\mathbb{Z}$ gives the $\infty$-category of bounded countable Tate objects, denoted $\mathscr{T}\mathrm{ate}_{{\aleph_{0}},R}^{\mathrm{b}}$.
  • Given an $\infty$-category $\mathscr{C}$ with finite limits, we let $\mathsf{Cond}(\mathscr{C})$ be the $\infty$-category of condensed objects in $\mathscr{C}$, defined as accessible $\mathscr{C}$-valued sheaves on profinite sets. For $\mathscr{C}=\mathscr{D}(R)$, note that there is an equivalence $\mathsf{Cond}(\mathscr{D}(R))\simeq\mathscr{D}(\mathsf{Cond}_{R})$ of stable $\infty$-categories.
Observe that the fully faithful functor $\mathsf{Mod}_{R}\hookrightarrow\mathsf{Cond}_{R}$ defined in (1.4) induces a fully faithful functor of stable $\infty$-categories $\mathscr{D}(R)\hookrightarrow\mathscr{D}(\mathsf{Cond}_{R})$, which is left adjoint to evaluation at $*\in\mathsf{ProFin}$, and factors through the stable full subcategory $\mathscr{D}(\mathsf{Solid}_{R})$ of solid objects. Beware that, as opposed to the previous sections, we will not distinguish between an object in $\mathscr{D}(R)$ and its image through this functor.
@@
Let $R$ be a ring of finite type over $\mathbb{Z}$. The restriction
$\mathscr{T}\mathrm{ate}_{\aleph_{0},R}^{\mathrm{b}}\longrightarrow\mathscr{D}(\mathsf{Cond}_{R})$
of the functor (3.1) to bounded countable Tate objects is fully faithful.
Proof.
First observe that the functor (3.1) factors through the full subcategory $\mathscr{D}(\mathsf{Solid}_{R})$, so that we may as well work in the latter $\infty$-category. Let $X$ and $Y$ be in $\mathscr{T}\mathrm{ate}_{{\aleph_{0}},R}^{\mathrm{b}}$. We want to prove that the comparison map
$\gamma_{X,Y}\colon\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{{\aleph_{0}},R}^{\mathrm{b}}}(X,Y)\longrightarrow\operatorname{Map}_{\mathscr{D}(\mathsf{Solid}_{R})}(\rho(X),\rho(Y))$ (3.2)
is an equivalence. By definition of Tate objects, we can pick elementary Tate objects $V$ and $W$ of which $X$ and $Y$ are respectively retracts. Replacing $V$ and $W$ by appropriate truncations if necessary, we may further assume that both are bounded. Observe that $\gamma_{X,Y}$ is a retract of $\gamma_{V,W}$; since equivalences are stable under retracts, it will suffice to show that $\gamma_{V,W}$ is an equivalence. As an elementary Tate object, $V$ fits in a fiber sequence $V^{\mathrm{p}}\to V\to V^{\mathrm{i}}$ with $V^{\mathrm{p}}\in\mathrm{Pro}_{\aleph_{0}}(\mathrm{Perf}_{R})$ and $V^{\mathrm{i}}\in\mathrm{Ind}_{\aleph_{0}}(\mathrm{Perf}_{R})$ by [Hen17b, Corollary 3.4]. Since $V$ is bounded, by applying appropriate truncations to the fiber sequence, we may assume without any loss of generality that $V^{\mathrm{p}}$ and $V^{\mathrm{i}}$ are also bounded pro and ind-objects. We decompose similarly $W$. Consider the induced commutative diagram
${{\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{R}}(V^{\mathrm{i}},W^{\mathrm{p}})}}$${{\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{R}}(V^{\mathrm{i}},W)}}$${{\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{R}}(V^{\mathrm{i}},W^{\mathrm{i}})}}$${{\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{R}}(V,W^{\mathrm{p}})}}$${{\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{R}}(V,W)}}$${{\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{R}}(V,W^{\mathrm{i}})}}$${{\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{R}}(V^{\mathrm{p}},W^{\mathrm{p}})}}$${{\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{R}}(V^{\mathrm{p}},W)}}$${{\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{R}}(V^{\mathrm{p}},W^{\mathrm{i}})}}$
with exact rows and columns, and the corresponding diagram for mapping spaces in $\mathscr{D}(\mathsf{Solid}_{R})$, we see that it suffices to show the result for $V$ and $W$ either in $\mathrm{Pro}^{\mathrm{b}}_{\aleph_{0}}(\mathrm{Perf}_{R})$ or in $\mathrm{Ind}^{\mathrm{b}}_{\aleph_{0}}(\mathrm{Perf}_{R})$. We treat the case where $V\in\mathrm{Pro}^{\mathrm{b}}_{\aleph_{0}}(\mathrm{Perf}_{R})$ and $W\in\mathrm{Ind}^{\mathrm{b}}_{\aleph_{0}}(\mathrm{Perf}_{R})\subseteq\mathscr{D}^{\mathrm{b}}(R)$, the other three cases being easier. Using that every object $V_{i}$ is compact, the proof reduces to the following claim. ∎
Lemma 3.3.
For $V\colon I\to\mathrm{Perf}_{R}$ in $\mathrm{Pro}^{\mathrm{b}}_{\aleph_{0}}(\mathrm{Perf}_{R})$ and $W\in\mathscr{D}^{\mathrm{b}}(R)$, the canonical map
$\operatorname*{colim}_{i}\operatorname{Map}_{\mathscr{D}(\mathsf{Solid}_{R})}(V_{i},W)\longrightarrow\operatorname{Map}_{\mathscr{D}(\mathsf{Solid}_{R})}(\lim_{i}V_{i},W)$ (3.3)
@@
Since $\mathsf{Solid}_{R}$ is an abelian category with enough projectives, we may consider the projective model structure on its category $\mathrm{Ch}^{-}(\mathsf{Solid}_{R})$ of right bounded chain complexes. Choose $a$ and $b$ such that each complex $V_{i}$ is of tor-amplitude in $[a,b-1]$. Using a fibrant replacement, the pro-object $V$ can be modelled by a diagram $I\to\mathrm{Ch}^{-}(\mathsf{Solid}_{R}),\,i\mapsto V_{i}$ in which each $V_{i}$ is a perfect complex of $R$-modules concentrated in degrees $[a,b]$ and each transition map $V_{i}\to V_{i-1}$ is an epimorphism.
Let us prove this last claim in details. Starting from any pro-object $i\mapsto V_{i}$ modelling $V$, we can inductively replace $f_{i}\colon V_{i}\to V_{i-1}$ by a map $f^{\prime}_{i}\colon V_{i}^{\prime}=V_{i}\oplus X_{i}\to V_{i-1}$ constructed as follows. Since $V_{i-1}$ is perfect, in every degree $p\in[a,b-1]$ we may choose $n_{p}\in\mathbb{N}$ and a surjection $g^{p}\colon R^{n_{p}}\to V_{i-1}^{p}$. Consider the standard contractible complex $D(k)=(\dots\to 0\to R\stackrel{{\scriptstyle\mathrm{id}}}{{\to}}R\to 0\to\dots)$ with $R$ in (cohomological) degrees $k$ and $k+1$. Observe that there is a unique morphism of complexes $\bar{g}^{p}\colon D(p)^{\oplus n_{p}}\to V_{i-1}$ that restricts to $g^{p}$ in degree $p$. Now take $X_{i}$ to be $\bigoplus_{p\in[a,b-1]}D(p)^{\oplus n_{p}}$ and $f^{\prime}_{i}$ to be the sum $f\oplus\bigoplus_{p\in[a,b-1]}\bar{g}^{p}$. By construction, $f_{i}\colon V_{i}\to V_{i-1}$ factors through the acyclic cofibration $V_{i}\stackrel{{\scriptstyle\sim}}{{\hookrightarrow}}V_{i}^{\prime}$, the map $f^{\prime}_{i}$ is degreewise surjective and $V_{i}^{\prime}$ is a perfect complex concentrated in degrees $[a,b]$, as desired.
It follows that the diagram $i\mapsto V_{i}$ we obtain is fibrant in the injective model structure on the category of diagrams $\operatorname{Fun}(I,\mathrm{Ch}^{-}(\mathsf{Solid}_{R}))$, so that there is a quasi-isomorphism $\lim_{i}V_{i}\simeq\operatorname*{holim}_{i}V_{i}$. We claim that the object $\lim_{i}V_{i}$ is actually cofibrant in the projective model structure. Since it is bounded, it suffices to show that it is degreewise projective. Now in degree $p$ the tower is given by surjections $V_{i}^{p}\to V_{i-1}^{p}$ between finitely generated projective modules, which then split, so that the limit $\lim_{i}V_{i}^{p}$ is a direct product $\prod_{i}Q_{i}^{p}$ of finitely generated projective modules $Q_{i}^{p}$. Since $R$ is a finitely generated $\mathbb{Z}$-algebra, the product $\prod_{\mathbb{N}}R$ is projective in $\mathsf{Solid}_{R}$. For every $i$ we may write $Q_{i}^{p}\oplus P_{i}^{p}\cong R^{m_{i,p}}$ for some modules $P_{i}^{p}$ and some $m_{i,p}\in\mathbb{N}$, so that we have $\prod_{i}Q_{i}^{p}\oplus\prod_{i}P_{i}^{p}\cong\prod_{i}R^{m_{i,p}}\cong\prod_{\mathbb{N}}R$ which implies that $\prod_{i}Q_{i}^{p}$ is also projective.
We have shown that $\lim_{i}V_{i}$ is cofibrant. As a consequence, there is an equivalence of Kan complexes
$\operatorname{Map}_{\mathscr{D}(\mathsf{Solid}_{R})}(\lim_{i}V_{i},W)\simeq\mathrm{DK}\mathopen{}\mathclose{{}\left(\tau^{\leqslant 0}\operatorname{\underline{Hom}}_{\mathrm{Ch}^{-}(\mathsf{Solid}_{R})}(\lim_{i}V_{i},W)}\right)$ (3.4)
where $\operatorname{\underline{Hom}}$ is the $\mathrm{Ch}(R)$-enriched Hom, $\tau^{\leqslant 0}\colon\mathrm{Ch}(R)\to\mathrm{Ch}(R)^{\leqslant 0}$ is the (canonical) truncation functor and
$\mathrm{DK}\colon\mathrm{Ch}(R)^{\leqslant 0}\simeq\mathsf{sMod}_{R}\to\mathsf{sSet}$
is the composite of the Dold–Kan equivalence with the forgetful functor. As every $V_{i}$ is also cofibrant, formula (4.2) also holds with $\lim_{i}V_{i}$ replaced by any of the $V_{i}$.
Now in every degree $p$ we have isomorphisms of $R$-modules
$\operatorname{\underline{Hom}}_{\mathrm{Ch}(\mathsf{Solid}_{R})}(\lim_{i}V_{i},W)^{p}$ $=\bigoplus_{q\in[a,b]}\operatorname{Hom}_{\mathsf{Solid}_{R}}(\lim_{i}V_{i}^{q},W^{p+q})$
$\cong\bigoplus_{q\in[a,b]}\operatorname*{colim}_{i}\operatorname{Hom}_{\mathsf{Solid}_{R}}(V_{i}^{q},W^{p+q})$
$\cong\operatorname*{colim}_{i}\bigoplus_{q\in[a,b]}\operatorname{Hom}_{\mathsf{Solid}_{R}}(V_{i}^{q},W^{p+q})$
$\cong\operatorname*{colim}_{i}\operatorname{\underline{Hom}}_{\mathrm{Ch}(\mathsf{Solid}_{R})}(V_{i},W)^{p}$
using the $1$-categorical embedding $\mathsf{Tate}_{{\aleph_{0}},R}\hookrightarrow\mathsf{Solid}_{R}$ (Proposition 1.5) in the second line. By naturality of these isomorphisms, we obtain an isomorphism of chain complexes
$\operatorname{\underline{Hom}}_{\mathrm{Ch}(\mathsf{Solid}_{R})}(\lim_{i}V_{i},W)\cong\operatorname*{colim}_{i}\operatorname{\underline{Hom}}_{\mathrm{Ch}(\mathsf{Solid}_{R})}(V_{i},W).$
Finally using that both functors $\mathrm{DK}$ and $\tau^{\leqslant 0}$ commute with filtered colimits, we deduce that
$\operatorname{Map}_{\mathscr{D}(\mathsf{Solid}_{R})}(\lim_{i}V_{i},W)$ $\simeq\mathrm{DK}\mathopen{}\mathclose{{}\left(\tau^{\leqslant 0}\operatorname{\underline{Hom}}_{\mathrm{Ch}(\mathsf{Solid}_{R})}(\lim_{i}V_{i},W)}\right)$
$\cong\mathrm{DK}\mathopen{}\mathclose{{}\left(\tau^{\leqslant 0}\operatorname*{colim}_{i}\operatorname{\underline{Hom}}_{\mathrm{Ch}(\mathsf{Solid}_{R})}(V_{i},W)}\right)$
$\cong\operatorname*{colim}_{i}\mathrm{DK}\mathopen{}\mathclose{{}\left(\tau^{\leqslant 0}\operatorname{\underline{Hom}}_{\mathrm{Ch}(\mathsf{Solid}_{R})}(V_{i},W)}\right)$
$\simeq\operatorname*{colim}_{i}\operatorname{Map}_{\mathscr{D}(\mathsf{Solid}_{R})}(V_{i},W),$
which concludes the proof. ∎